You are standing on the platform in a high speed train station. The train is passing you and is emitting a
frequency of 560 HZ. If the temperature of the surrounding air is 295.2622 kelvin, what must the speed of
the train be if the Doppler frequency is only of the source?
1
Expert's answer
2019-01-16T10:28:11-0500

Calculate the apparent frequency of the train by the relation as follows:

c/f'=c/f+V_s/f


Here, the velocity of the source isV_s , frequency of source is f and the velocity of sound is c.

Calculate the velocity of sound by the relation as follows:

C = 331.5+0.607 T

Here T is the temperature.

Substitute T =22.2622 celcius

C= 331.5+0.607(22.2622 celcius) =334.15 m/s

Then substitute the values in the formula

334.15/560=334.15/560+V_s/560

V_s=560 m/s